OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (anonymous):

( negative 1, and negative 7)

OpenStudy (agl202):

x1 = - 2 y1 = 3 x2 = - 5 y2 = 6 Slope, m = (y2 - y1) / (x2 - x1) m = (6 - 3) / [- 5 - (- 2)] m = 3 / (- 5 + 2) m = 3 / - 3 m = - 1 Hope this Helped!!! =)

OpenStudy (anonymous):

Thanks! :)

OpenStudy (agl202):

Ur Welcome!!!